# section[1.5em]1.5em1pc

### v2

@@ 14.   Associative algebras versus restricted Lie algebras
Therefore,
$\eta_{L}^{*}(-)_{\operatorname{RLie}}(u(L)\ltimes M\to u(L))=\eta_{L}^{*}((u(L)\ltimes M)_{\operatorname{RLie}}\to u(L)_{\operatorname{RLie}})=L\times_{u(L)_{\operatorname{RLie}}}(u(L)\ltimes M)_{\operatorname{RLie}}\to L,$
and $\phi\colon L\times_{u(L)_{\operatorname{RLie}}}(u(L)\ltimes M)_{\operatorname{RLie}}\to L$ is an abelian group object in $(\operatorname{RLie}/L)_{\mathrm{ab}}$. The action of $L$ on $M:=\ker\phi$ is given by
@@
15.   Good triple of operads and comparison maps in Quillen (co)homology
J.-L. Loday defined and studied generalized bialgebras and triple of operads in [Loday08]. In this section we prove a comparison isomorphism theorem for Quillen cohomology in the context of good triple of operads. Only in this section we suppose that the ground field $\mathbb{F}$ has characteristic zero. In that case, if $\mathcal{P}$ is an operad then the norm map $\mathrm{Tr}\colon S(\mathcal{P})\rightarrow\Gamma(\mathcal{P})$ is a natural isomorphism [Fresse00]. Therefore a $\Gamma(\mathcal{P})$-algebra doesn’t carry more structure than a $\mathcal{P}$-algebra.
@@ 15.   Good triple of operads and comparison maps in Quillen (co)homology
Let $(\mathcal{C},\mathcal{A},\mathcal{P})$ be a good triple of operads. Then $\mathcal{P}$ is the primitive operad $\mathrm{Prim}_{\mathcal{C}}\,\mathcal{A}$, which is a suboperad of $\mathcal{A}$. The inclusion $\mathrm{Prim}_{\mathcal{C}}\,\mathcal{A}\subseteq\mathcal{A}$ induces a functor
@@
which is a forgetful functor in a sense that the composition
@@
is the forgetful functor $\mathcal{A}\text{-}\mathrm{Alg}\to\mathrm{Vect}_{\mathbb{F}}$. The functor $G$ admits a left adjoint functor
@@
$E_{0}(U(L))\cong\mathcal{Z}(L)$
and thus
$E_{1}(U(L))\cong H_{*}(\mathcal{Z}(L)).$
Let $L,L^{\prime}$ be two simplicial $\mathcal{P}$-algebras and $f\colon L\to L^{\prime}$ a weak equivalence. The morphism $U(f)\colon U(L)\to U(L^{\prime})$ preserves filtrations and it induces a morphism of spectral sequences
$E_{r}(U(L))\to E_{r}^{\prime}(U(L^{\prime})).$
Since $f$ is a weak equivalence, it follows by a theorem of Dold [Dold58] that
@@
Let $\mathcal{T}$ be a operad and $S$ a $\mathcal{T}$-algebra. The category $S$-modules over $\mathcal{T}$ is equivalent to the category of abelian group objects of $\mathcal{T}\text{-}\mathrm{Alg}/S$ i.e the category of Beck $S$-modules in the category of $\mathcal{T}$-algebras (see [LodayV12], [GoerssH00]). Hence the category of $U(P)$-modules over the operad $\mathcal{A}$ is equivalent to the category of Beck $U(P)$-modules. Under this equivalence the $U(P)$-module $M$ is associated to the abelian group object $U(P)\ltimes M\to U(P)$, where by $U(P)\ltimes M$ we denote the semidirect product of $U(P)$ by $M$ in the category of $\mathcal{A}$-algebras. We have the following pullback diagram in the category of $\mathcal{P}$-algebras,
@@
If we consider the good triple $(\mathrm{Com},\operatorname{As},\operatorname{Lie})$ then by 15.4 we recover in characteristic zero a well known result [CartanE56]*§XIII, Theorem 5.1. In particular, let $L$ be a Lie algebra over $\mathbb{F}$ and $M$ a $U(L)$-bimodule. By 15.4 we have the following isomorphism
@@
where ${}_{L}M$ is $M$ viewed as a left Lie $L$-module via the action $l\cdot m:=lm-ml$ for all $l\in L$ and $m\in M$. Moreover, Quillen cohomology for the category of associative algebras is shifted Hochschild cohomology of associative algebras and Quillen cohomology of Lie algebras is shifted Chevalley–Eilenberg cohomology of Lie algebras (see [Barr96]). In other words we have
@@
In [Loday01] J.-L. Loday defined the notion of dendriform algebra which dichotomizes the notion of associative algebra.
@@
M. Ronco in [Ronco02] proves a Milnor–Moore type theorem for dendriform algebras. In particular, there is a good triple of operads $(\operatorname{As},\mathcal{D},\mathcal{B})$. The concept of bimodule over a dendriform algebra is defined by M. Aguiar in [Aguiar04]. Cohomology of dendriform algebras with coefficients in bimodules has been studied by A. Das in [Das22].
Let $D$ be a dendriform algebra. The notion of Beck $D$-module is equivalent to the notion of bimodule over $D$. By 15.4 we have the following.
@@
Let $B$ be a $\mathcal{B}$-algebra and $M$ a $U_{dend}(B)$-bimodule, where $U_{dend}(B)$ denotes the enveloping dendriform algebra of the brace algebra $B$. Then we have the following isomorphism
@@
In positive characteristic, for classical bialgebras we proved 14.2. We notice that the primitives of a classical bialgebra form a restricted Lie algebra, i.e., a $\Gamma(\operatorname{Lie})$-algebra. If one wants to extend this result to the context of generalized bialgebras, the right framework seems to be the category of $\Gamma(\mathcal{P})$-algebras.

### v3

@@ 14.   Associative algebras versus restricted Lie algebras
Applying the functor $\eta_{L}^{*}(-)_{\operatorname{RLie}}$ to the map $u(L)\ltimes M\to u(L)$ yields a map
$\phi\colon L\times_{u(L)_{\operatorname{RLie}}}(u(L)\ltimes M)_{\operatorname{RLie}}\to L$
which is an abelian group object in $(\operatorname{RLie}/L)_{\mathrm{ab}}$. The action of $L$ on $M:=\ker\phi$ is given by
@@
15.   Good triples of operads and comparison maps in Quillen (co)homology
J.-L. Loday defined and studied generalized bialgebras and triples of operads in [Loday08]. In this section we prove a comparison isomorphism theorem for Quillen cohomology in the context of good triples of operads. Only in this section we suppose that the ground field $\mathbb{F}$ has characteristic zero. In that case, if $\mathcal{P}$ is an operad then the norm map $\mathrm{Tr}\colon S(\mathcal{P})\rightarrow\Gamma(\mathcal{P})$ is a natural isomorphism [Fresse00]. Therefore a $\Gamma(\mathcal{P})$-algebra doesn’t carry more structure than a $\mathcal{P}$-algebra.
@@ 15.   Good triples of operads and comparison maps in Quillen (co)homology
Let $(\mathcal{C},\mathcal{A},\mathcal{P})$ be a good triple of operads, as defined in [Loday08]*2.4.1, 2.5.6. Then $\mathcal{P}$ is the primitive operad $\mathrm{Prim}_{\mathcal{C}}\,\mathcal{A}$, which is a suboperad of $\mathcal{A}$. The inclusion $\mathrm{Prim}_{\mathcal{C}}\,\mathcal{A}\subseteq\mathcal{A}$ induces a functor
@@
$\begin{cases}E_{0}(U(L))\cong\mathcal{Z}(L)\\ E_{1}(U(L))\cong H_{*}(\mathcal{Z}(L)).\end{cases}$
Let $L,L^{\prime}$ be two simplicial $\mathcal{P}$-algebras and $f\colon L\to L^{\prime}$ a weak equivalence. The morphism $U(f)\colon U(L)\to U(L^{\prime})$ preserves filtrations and it induces a morphism of spectral sequences $E_{r}(U(L))\to E_{r}(U(L^{\prime}))$. Since $f$ is a weak equivalence, it follows by a theorem of Dold [Dold58] that
@@
Let $\mathcal{T}$ be a operad and $S$ a $\mathcal{T}$-algebra. The category of $S$-modules over $\mathcal{T}$ is equivalent to the category of abelian group objects of $\mathcal{T}\text{-}\mathrm{Alg}/S$, i.e., the category of Beck $S$-modules in the category of $\mathcal{T}$-algebras (see [LodayV12, GoerssH00]). Hence the category of $U(P)$-modules over the operad $\mathcal{A}$ is equivalent to the category of Beck $U(P)$-modules. Under this equivalence the $U(P)$-module $M$ is associated to the abelian group object $U(P)\ltimes M\to U(P)$, where by $U(P)\ltimes M$ we denote the semidirect product of $U(P)$ by $M$ in the category of $\mathcal{A}$-algebras. We have the following pullback diagram in the category of $\mathcal{P}$-algebras,
@@
Example 15.5.
If we consider the good triple $(\mathrm{Com},\operatorname{As},\operatorname{Lie})$, then by Theorem 15.4 we recover in characteristic zero a well known result [CartanE56]*§XIII, Theorem 5.1. In particular, let $L$ be a Lie algebra over $\mathbb{F}$ and $M$ a $U(L)$-bimodule. By Theorem 15.4 we have the following isomorphism
@@
where ${}_{L}M$ is $M$ viewed as a left Lie $L$-module via the action $l\cdot m:=lm-ml$ for all $l\in L$ and $m\in M$. Moreover, Quillen cohomology for the category of associative algebras is shifted Hochschild cohomology of associative algebras and Quillen cohomology of Lie algebras is shifted Chevalley–Eilenberg cohomology of Lie algebras (see [Barr96]). In other words, we have
@@
As another example, let us consider dendriform algebras, defined by J.-L. Loday in [Loday01].
@@
M. Ronco in [Ronco02] proved a Milnor–Moore type theorem for dendriform algebras. In particular, there is a good triple of operads $(\operatorname{As},\mathcal{D},\mathcal{B})$. The concept of bimodule over a dendriform algebra was defined by M. Aguiar in [Aguiar04]. Cohomology of dendriform algebras with coefficients in bimodules has been studied by A. Das in [Das22].
Let $D$ be a dendriform algebra. The notion of Beck $D$-module is equivalent to the notion of bimodule over $D$. Applying Theorem 15.4 to the good triple $(\operatorname{As},\mathcal{D},\mathcal{B})$ yields the following.
@@
Let $B$ be a $\mathcal{B}$-algebra, and let $M$ be a $U_{dend}(B)$-bimodule, where $U_{dend}(B)$ denotes the enveloping dendriform algebra of the brace algebra $B$. We have the following isomorphism:
@@
In positive characteristic, for classical bialgebras we proved Theorem 14.2. We notice that the primitives of a classical bialgebra form a restricted Lie algebra, i.e., a $\Gamma(\operatorname{Lie})$-algebra. If one wants to extend this result to the context of generalized bialgebras, the right framework seems to be the category of $\Gamma(\mathcal{P})$-algebras.